OpenStudy (anonymous):

Please help me solve story problem: In a certain city, it was recorded that 15 1/5 inches of rain fell over 24 hour period, how much rain fell per hour on average?

OpenStudy (anonymous):

rate of rain *hours = total rain. r * 24 = 15.2 r = 15.2 / 24 r = 0.633

OpenStudy (ybarrap):

15.20 inches per 24 hour period => (15.2/24) in/hour - That's it

OpenStudy (anonymous):

thanks 19/30 was the correct answer